The pressure or air in a reasonably good vacuum might be 2.0 x 10-5 mmHg. What mass of air exists in a 250mL volume at this pressure and 250C? Take M=28 Kg/kmol for air.


1
Expert's answer
2021-04-20T16:26:14-0400

Let's assume that air in the given conditions can be an ideal gas. Then

"\displaystyle pV = \frac{m}{M} RT"

Converting everything to SI units

"p =2.0 \cdot 10^{-15} \cdot 133.32 = 266.64 \cdot 10^{-15}\; Pa"

"V= 250 \cdot 10^{-6} \; m^3"

"T = 25 + 273 =298\; K"

"M = 28 \cdot 10^{-3}\; kg/ mol"


"\displaystyle m = \frac{pVM}{RT} = \frac{266.64 \cdot 10^{-15} \cdot 250 \cdot 10^{-6} \cdot 28 \cdot 10^{-3}}{8.31 \cdot 298} =753.7 \cdot 10^{-24} \; kg = 7. 537 \cdot 10^{-22} \;kg"

Answer: "7. 537 \cdot 10^{-22} \;kg"